OpenStudy (anonymous):

please help i dont get this 11/13 multiply by what equals 1

OpenStudy (anonymous):

11/13= .8461 .8461 X = 1 x = 1/.8461 X= 1.1818

OpenStudy (gw2011):

(11/13)(13/11)=143/143=1 Another way is to cancel: the numerator and denominator of 11 cancel to 1 and the numerator and denominator of 13 cancel to 1. Therefore, you get 1/1=1